# Numerical simulation of type II primordial black hole formation

### v2

@@ 4.1 PBH mass with the Gaussian initial fluctuation
Here, we evaluate the mass of the PBH in two different ways at each characteristic time. The one is to calculate the Misner-Sharp mass inside the apparent horizon at its first appearance with the future-trapped region (see Fig. 6). We refer to the mass at this time as the initial black hole mass. The apparent horizon appears as a future outer trapping horizon in a type I PBH, while it appears as a bifurcating outer trapping horizon in type II. We also calculate the final PBH mass by estimating the effect of the accretion. It is known that the accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [2, 42]. In this accretion model, the time evolution of the mass is given by
@@
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
The initial profiles of $\zeta_{\rm cFLRW}$, $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 12, 13 and 14. This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 13: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$. By taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
@@ 5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH and a type II fluctuation does in a type II PBH. However, in the radiation fluid case, this is not true. We found a gap between the classifications in terms of the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs in terms of the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [5], the spacetime of PBH is described by the patchwork of the three different regions : the innermost overdense region given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 16, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;*)(+;*)( + ; * ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH horizon and cosmological horizon become future outer (;)(-;-)( - ; - ) and past inner (+;+)(+;+)( + ; + ), denoted by red null and blue timelike lines, respectively.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 16, and the resultant spacetime diagram is shown in the left panel of Fig. 18. In these conformal diagrams, we define the right and left directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
For the type II fluctuation, the extracted regions are yellow-shaded in Fig. 17. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for the type II is given by the right panel of Fig. 18.
(a) closed FLRW for dust
(b) Schwarzschild
@@
Figure 16: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 18 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past outer (+;)(+;-)( + ; - ), future inner (;+)(-;+)( - ; + ), and past inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red null, blue null, red timelike, and blue timelike lines, respectively.
(a) closed FLRW for dust
(b) Schwarzschild
@@
Figure 18: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I PBH as a causal structure of typical case PBH. The right one is constructed by changing the cutting diagram range and defining a type II PBH.
(a) closed FLRW for w=1/3
@@
Figure 19: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null.

### v3

@@ 4.1 PBH mass with the Gaussian initial fluctuation
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref. [1].
@@
The initial profiles of $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 12 and 13. The tail profile is different from the original three-zone model, but it is similar to the three-zone profile around the peak of $\zeta$. This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 12: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$. Taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$. The original profile near the peak $\zeta_{\rm cFLRW}$ [1] is plotted in Fig. 19 in Appendix B.
@@ 5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH. However, in the case of radiation fluids, this is not true. We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 15, and the resultant spacetime diagram is shown in the left panel of Fig. 17. These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
The extracted regions are yellow-shaded for the type II fluctuation in Fig. 16. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for type II is given by the right panel of Fig. 17.
@@
Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 17 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue -solid, red -dashed, and blue -dashed lines, respectively.
@@
Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I PBH as a typical case PBH causal structure. The right one is constructed by changing the cutting diagram range and defining a type II PBH.
@@
Figure 18: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null. Future-degenerate $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;0)$ and past-degenerate $(+;0)$ trapping horizons are denoted by red-dotted and blue-dotted lines, respectively.
@@ Appendix B Derivation of $\zeta_{\text{cFLRW}}$ in radial coordinate $r$
This functional form is shown as:
Appendix C Double-peaked compaction function and averaged density perturbation in the type II fluctuation
For type II fluctuations, let us assume that $R^{\prime}=\partial_{r}R$, where $R\coloneqq are^{\zeta}$ is the areal radius, satisfies the following condition
$R^{\prime}\begin{cases}>0&(0<r<r_{m1},r_{m2}<r)\\ =0&(r=r_{m1},r_{m2})\\ <0&(r_{m1}<r<r_{m2})\end{cases}.$ (C.1)
Since
$R^{\prime}=ae^{\zeta}(1+r\zeta^{\prime}),$ (C.2)
we find $r\zeta^{\prime}>-1$ for $0<r<r_{m1}$, $r_{m2}<r$, $r\zeta^{\prime}=-1$ for $r=r_{m1},r_{m2}$, and $r\zeta^{\prime}<-1$ for $r_{m1}<r<r_{m2}$. From the above assumption, we can deduce that $r\zeta^{\prime}$ must have at least one minimum below $-1$ with $(r\zeta^{\prime})^{\prime}=0$ between $r_{m1}$ and $r_{m2}$. For simplicity, we additionally assume within this section that $r\zeta^{\prime}$ takes only one minimum at $r=r_{m3}$ between $r_{m1}$ and $r_{m2}$ and $r\zeta^{\prime}$ monotonically decreases for $0<r<r_{m3}$ and monotonically increases for $r_{m3}<r$. Since we have the compaction function 777 One can rewrite $C_{\rm SS}(r)$ as $C_{\rm SS}(r)=-r\zeta^{\prime}\quantity(1+\frac{1}{2}r\zeta^{\prime})=\frac{3}{4}\delta_{\rm l}\quantity(1-\frac{3}{8}\delta_{\rm l})=-\frac{9}{32}\left(\delta_{l}-\frac{4}{3}\right)^{2}+\frac{1}{2},$ where $\delta_{\rm l}\coloneqq-\frac{4}{3}r\zeta^{\prime}.$ is the ”linear-order density perturbation”. Therefore, $C_{\rm SS}$ as a function of $\delta_{\rm l}$ takes two zeroes at $\delta_{\rm l}=0$ and $\delta_{\rm l}=8/3$ and a maximum $1/2$ at $\delta_{\rm l}=4/3$. The type II fluctuation is characterized by the existence of the interval in $r$ with $\delta_{l}(r)>4/3$, where $C_{\rm SS}(r)$ as a function of $r$ takes two maxima $1/2$ at its boundaries with $\delta_{l}(r)=4/3$.
$\mathcal{C}_{\rm SS}(r)=\frac{1}{2}\quantity(1-\quantity(1+r\zeta^{\prime})^{2})$ (C.3)
and its derivative
$\mathcal{C}_{\rm SS}^{\prime}(r)=-\quantity(1+r\zeta^{\prime})\quantity(\zeta^{\prime}+r\zeta^{\prime\prime}),$ (C.4)
we find that the compaction function $\mathcal{C}_{\rm SS}(r)$ has three extrema at $r=r_{m1}$, $r_{m3}$ and $r_{m2}$ with $\mathcal{C}_{\rm SS}^{\prime}(r)=0$, two of which at $r=r_{m1}$ and $r_{m2}$ correspond to two peaks with the value $1/2$, and the rest at $r=r_{m3}$ to its minimum below $1/2$. This feature is typically seen for type II fluctuations and actually the case for the Gaussian-shaped profile Eq. (2.5) as seen in Fig. 20.
Refer to captionRefer to caption
Figure 20: The situation of type II fluctuation $r\zeta^{\prime}<-1$ from non-monotonicity of the areal radius $R$, $R^{\prime}<0$, and corresponding to the two peaks of the compaction function $\mathcal{C}_{\rm SS}$ (or averaged density perturbation $\bar{\delta}$ at the horizon entry).
The density perturbation $\delta\coloneqq\rho/\rho_{\rm b}-1$ in a sub-leading order ($\order{\epsilon^{2}}$ ) of the long-wavelength solutions for uniform Hubble (constant mean curvature) slice is written as [21],
$\delta\simeq-\frac{4}{3}e^{-5\zeta/2}\bar{\Delta}\quantity(e^{\zeta/2})\quantity(\frac{1}{aH_{\rm b}})^{2},$ (C.5)
where $\bar{\Delta}$ is the Laplacian of the flat background geometry. The compaction function can be identified with the volume averaged density perturbation at the horizon entry $R=1/H_{\rm b}$,
$\left.\bar{\delta}\right|_{H_{\rm b}=1/R}\simeq 2\mathcal{C}_{\rm SS}.$ (C.6)
Then, the plot of $\mathcal{C}_{\rm SS}$ in Fig. 20 can also be understood as $\bar{\delta}$ up to the factor of $2$. For the type II fluctuations ($\mu\gtrsim 1.4$), the value of $\bar{\delta}$ takes a minimal value at $r=\sqrt{2}/k\simeq 0.14$ with $k=10$, and the minimal value decreases with increasing $\mu$.
It is worth noting that the separate universe (pinch-off) limit will not occur for this Gaussian-shaped profile of $\zeta$, while Kopp et al. considered the limit based on the other profile via the closed FLRW (or Three-zone) geometry of Eq. (4.3) (or Eq.(4.4)).